# Requirements Coverage-Guided Minimization for Natural Language Test Cases

### v2

@@ Abstract.
As software systems evolve, test suites tend to grow in size and often contain redundant test cases. Such redundancy increases testing effort, time, and cost. Test suite minimization (TSM) aims to eliminate such redundancy while preserving key properties, such as requirement coverage and fault-detection capability. In this paper, we propose RTM (Requirement coverage-guided Test suite Minimization), a novel TSM approach designed for requirement-based testing (validation), which can effectively reduce test suite redundancy while ensuring full requirement coverage and a high fault detection rate (FDR) under a fixed minimization budget. Based on common practice in critical systems where functional safety is important, we assume that test cases are specified in natural language and traced to requirements before implementation. RTM utilizes text embedding technique to convert test cases into vector representations, on which a distance function is employed to compute similarity values between test case pairs. Guided by these similarity values, a Genetic Algorithm (GA) whose population is initialized using a coverage-preserving strategy is then employed to search for an optimal subset of diverse test cases that matches the budget. We investigate three preprocessing methods for test cases, seven different text embedding techniques, three distance functions, and three initialization strategies for the GA. We evaluate RTM on an industrial automotive system dataset comprising $736$ system test cases and $54$ requirements. Experimental results show that, while being scalable in terms of runtime, RTM outperforms all the baseline techniques in terms of FDR on most minimization budgets while maintaining full requirement coverage. Furthermore, we investigate the impact of test suite redundancy levels on the effectiveness of TSM, providing new insights into optimizing requirement-based test suites under practical constraints.
Test Suite Minimization, Requirements Testing, Genetic Algorithm, Automotive System
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: xxx; 2025; XXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Requirements analysis
@@ 1. Introduction
Software testing is a critical activity throughout the software development life cycle, playing a key role in early defect detection and in ensuring requirements compliance (myers2011art; everett2007software). In particular, requirements-driven testing is conducted to validate systems before deployment (dustin2009implementing). This is especially critical for systems that must demonstrate functional safety, such as automotive systems. In such contexts, it is also common to specify test cases in natural language (following templates) to facilitate their analysis and traceability to requirements, before manually writing the corresponding test code.
@@
However, due to the increasing complexity of systems and their continuous change by large teams, test suites are hard to maintain, tend to grow, and are often redundant (i.e., some of the test cases are likely to detect the same faults), thus leading to waste of time and resources (rothermel2002empirical). Further, in practice, testers have limited resources and time to test new system updates. To address this, given a test budget, test suite minimization (TSM) is proposed to prune test cases that are most likely to be redundant, while satisfying coverage criteria and maximizing fault detection (harrold1993methodology; von1999test), thus significantly reducing testing cost and time while limiting the impact on system quality or compliance.
@@
While several approaches have been proposed for TSM (noemmer2019evaluation; miranda2017scope; coviello2018clustering; liu2011user; viggiato2022identifying; cruciani2019scalable; zhang2019uncertainty; hemmati2013achieving; pan2023atm; DBLP:journals/tse/PanGB24; marchetto2017combining; xia2021test; anwar2019hybrid; yoo2010using), most existing techniques rely on code and structural coverage criteria (e.g., statement or branch coverage). These techniques do not address the unique challenges of minimizing test suites driven by system requirements, with test cases often specified in natural language. Moreover, to the best of our knowledge, no existing solution can simultaneously maintain full requirement coverage while adhering to a fixed minimization budget, a common situation in industrial practice due to limited time and resources.
@@
To bridge this gap, we propose RTM (Requirements coverage-guided Test suite Minimization), a novel approach specifically designed for requirement-based testing. RTM aims to reduce test suite redundancy (in terms of fault detection) while ensuring full requirement coverage, making it particularly well-suited to the practical needs of our industrial partners and to critical systems that must demonstrate functional safety in general. RTM takes requirement-based test cases–composed of natural language test steps–as input and computes pairwise similarity between them. Specifically, we investigate three different preprocessing methods to normalize the test steps, followed by seven distinct text embedding techniques–both sentence-level and word-level–to transform the textual data into vector representations: Term Frequency–Inverse Document Frequency (TF-IDF), Universal Sentence Encoder (USE) (DBLP:journals/corr/abs-1803-11175), LongT5 (DBLP:conf/naacl/GuoAUONSY22), Amazon Titan Text Embedding V2 (aws2024titantv2), Word2Vec (DBLP:journals/corr/abs-1301-3781), GloVe (DBLP:conf/emnlp/PenningtonSM14), FastText (DBLP:journals/tacl/BojanowskiGJM17). We then measure the similarity between test cases using three different distance metrics: cosine similarity and Euclidean distance for sentence-level embeddings, and Word Mover’s Distance (WMD) (DBLP:conf/icml/KusnerSKW15) for word-level embeddings. Based on the computed similarities, we employ a Genetic Algorithm (GA) (Luke2013Metaheuristics) to search for the optimal test suite that minimizes similarity, adheres to a fixed budget, and ensures full requirement coverage. To help the GA efficiently find valid subsets, we devise three initialization strategies that generate an initial population of subsets satisfying both coverage and budget constraints.
@@
RTM was evaluated on an industrial automotive dataset consisting of $736$ system test cases across seven test runs, along with $54$ requirements. The results demonstrate that RTM outperforms the baselines in terms of fault detection rate (FDR) under various minimization budgets. Moreover, we explore the impact of test suite redundancy on FDR. Our findings show that: ① RTM consistently outperforms all baseline techniques in terms of FDR across different test suite redundancy levels; ② The level of redundancy has a strong effect on the achieved FDR, for all TSM techniques; ③ Especially at lower minimization budgets and for highly redundant test suites, there is still significant room for improvement between the theoretical maximum FDR and that of RTM. ④ The runtime of RTM scales approximately linearly with test suite size, with small (i.e., of no practical consequence) absolute differences compared to the baselines. Given its consistently higher FDR performance across various test suites with different redundancy levels, RTM is therefore a better choice in many practical contexts.
@@
  • To foster reproducibility and facilitate further research, we have open-sourced our replication package, including source code, input data, and results for all research questions that allow the reproducibility of the experiments.
@@ 2.1. Test Cases based on Automotive Requirements
Like other critical systems, modern automotive software systems are expected to comply with functional safety standards such as ISO 26262 (siegl2010model). To ensure compliance, system engineers derive and maintain extensive test suites based on system requirements, which are executed to validate system behavior against expected requirements. Typically, test cases are specified in natural language, following templates, and then manually converted into test code and implemented by experts.
@@
2.2. Test Suite Minimization
Test suite minimization (TSM) aims to reduce the test suite size while ensuring key testing objectives, such as requirement coverage, fault-detection capability, and test-case diversity (harrold1993methodology; von1999test). Given that test suite diversity is reported to be positively correlated with fault detection capability (hemmati2013achieving), minimizing test case similarity can enhance the fault detection capability of the resulting test suite. When appropriately applied, TSM can significantly reduce testing costs and time while minimizing its impact on system quality or compliance.
Let $\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ be the set of system requirements for an automotive system, and let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{m}\}$ be the corresponding test suite. Each test case $t_{i}\in\mathcal{T}$ covers a subset of requirements denoted by $Cover(t_{i})\subseteq\mathcal{R}$. Note that in our dataset, each test case only covers one requirement, whereas many test cases can cover each requirement. This is because each test case is specifically designed to validate a scenario in which the requirement is expected to hold. Therefore, in our context, guaranteeing full requirement coverage requires at least as many test cases as requirements in the test suites. Moreover, covering all requirements is a must in automotive testing, to comply with functional safety standards (siegl2010model), which are adhered to by our industry partner. Let $Sim(t_{i},t_{j}):\mathcal{T}\times\mathcal{T}\rightarrow[0,1]$ denote the similarity between test cases $t_{i}$ and $t_{j}$.
@@ 3. Approach
This section introduces RTM, a TSM approach that satisfies the three criteria outlined in Section 2.2. It targets test cases derived from system requirements for validation. In short, it leverages a text embedding technique combined with a distance function to compute test case similarity, then employs evolutionary search to minimize the test suite. The primary objective of RTM is to find the optimal $\mathcal{T}^{\prime}$ that maximizes fault detection capability by minimizing the similarity values between the test cases while reducing the testing cost, measured in terms of test suite size, as well as ensuring $100\%$ requirement coverage.
@@
Figure 2. Overall Framework of RTM.:
@@ 3.3. Similarity Measure
For word-level embeddings (i.e., Word2Vec, GloVe, and FastText), we use WMD (DBLP:conf/icml/KusnerSKW15), which formulates the similarity between two test cases as an optimal transport problem. Specifically, WMD computes the minimum cumulative cost required to move the probability mass of all the words in one test case to those in another in the embedding space. The WMD is defined as:
@@
This optimization problem is equivalent to the Earth Mover’s Distance (monge1781memoire; wolsey1999integer; rubner1998metric) and can be solved via linear programming, with a computational complexity of approximately $O(n^{3}\log n)$ for $n$ unique words. Although the computation can be expensive when there are many words, WMD is well-suited to word-level embeddings because pairwise distances between the word embeddings approximate semantic dissimilarity, allowing the transport cost to reflect meaning differences between test cases (DBLP:conf/icml/KusnerSKW15).
In order to bound the WMD between 0 and $1$, and convert it from a distance function to a similarity measure, we normalize it as follows:
@@
For sentence-level embeddings (i.e., TF-IDF, USE, LongT5, and Amazon Titan Embedding V2), we investigate both cosine similarity and Euclidean distance, which have been widely used in machine learning, information retrieval, and NLP tasks to measure similarity or dissimilarity between vectors (gomaa2013survey; huang2008similarity; li2013distance; salton1988term). However, they operate differently and are suited to different use cases: cosine similarity measures the angle between two vectors, ignoring magnitude, whereas Euclidean distance measures the straight-line distance between two vectors. By applying different types of similarity metrics, we aim to explore how different vector representations and distance functions affect the ability of RTM to detect redundant test cases.
@@ 3.4. Evolutionary Search for Test Suite Minimization
As discussed in Section 2.2, the TSM in our context is an NP-hard problem. Therefore, to efficiently find a near-optimal, feasible solution, we apply a GA–the most widely used evolutionary search algorithm in software testing and the approach adopted by the state-of-the-art (SOTA) TSM work (i.e., ATM (pan2023atm) and LTM (DBLP:journals/tse/PanGB24))–which has demonstrated high effectiveness and efficiency for TSM. To tailor GA to our problem, we define the individuals (minimized test suites), the initialization strategy, selection, crossover, and mutation operators, and the fitness function that guides the search. The details are given below.
@@ 3.4.3. Search Process
After initializing a population with valid individuals (subsets), GA (golberg1989genetic) iteratively evolves this population using selection, crossover, and mutation operators towards better individuals with lower fitness values. In each generation, offspring are created by recombining and mutating the genes (i.e., selected test cases) of existing individuals. Each individual is evaluated by computing its fitness, and those with lower fitness are selected for the next generation. The evolutionary process is repeated until convergence, defined as a decrease in the fitness score of less than $0.0025$ across generations. The final output is the best subset identified during the search, which satisfies all constraints and minimizes the redundancy among selected test cases, indicated by the lowest fitness value.
@@
  • Crossover: We used the customized crossover operator (pymoo) that ensures the offspring maintains a fixed subset size. It first includes all test cases shared by both parent subsets, then adds test cases from either parent subset until the minimization budget is met.
  • Mutation: We used an inversion mutation operator (pymoo) that randomly selects a segment of the subset and reverses its order. Since a subset is represented as a binary vector, this approach preserves the subset size.
@@
We follow the published guidelines  [^1] https://www.obitko.com/tutorials/genetic-algorithms/recommendations.php used in prior work (DBLP:journals/tse/PanGB24) for setting GA hyperparameters. Specifically, we use a population size of $100$, a mutation rate of $0.01$, and a crossover rate of $0.90$. For each generation, the size of each solution (i.e., subset) remains fixed at the minimization budget.
@@ 2. RQ2:
    This RQ aims to compare the performance of RTM with two baseline approaches, namely Random Minimization and FAST-R (cruciani2019scalable), under various minimization budgets (i.e., 10%, 20%, …, 90%).

### v3

@@ Abstract.
As software systems evolve, test suites tend to grow in size and often contain redundant test cases. Such redundancy increases testing effort, time, and cost. Test suite minimization (TSM) aims to eliminate such redundancy while preserving key properties, such as requirement coverage and fault-detection capability. In this paper, we propose RTM (Requirement coverage-guided Test suite Minimization), a novel TSM approach designed for requirement-based testing (validation), which can effectively reduce test suite redundancy while ensuring full requirement coverage and a high fault detection rate (FDR) under a fixed minimization budget. Based on common practice in critical systems where functional safety is important, we assume that test cases are specified in natural language and traced to requirements before implementation. RTM utilizes text embedding technique to convert test cases into vector representations, on which a distance function is employed to compute similarity values between test case pairs. Guided by these similarity values, a Genetic Algorithm (GA) whose population is initialized using a coverage-preserving strategy is then employed to search for an optimal subset of diverse test cases that matches the budget. We investigate three preprocessing methods for test cases, seven different text embedding techniques, three distance functions, and three initialization strategies for the GA. We evaluate RTM on an industrial automotive system dataset comprising $736$ system test cases covering $54$ requirements. Experimental results show that, while being scalable in terms of runtime, RTM outperforms all the baseline techniques in terms of FDR on most minimization budgets while maintaining full requirement coverage. Furthermore, we investigate the impact of test suite redundancy levels on the effectiveness of TSM, providing new insights into optimizing requirement-based test suites under practical constraints.
Test suite minimization, Test suite reduction, Requirements testing, Natural language processing, Large language models, Genetic algorithm, Automotive system
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: xxx; 2025; XXXXisbn: 978-1-4503-XXXX-X/2018/06copyright: ccjournal: TOSEMjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3798164ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Requirements analysis
[…]
by-nc-nd
@@ 1. Introduction
Software testing is a critical activity throughout the software development life cycle, playing a key role in early defect detection and in ensuring requirements compliance (Myers et al., 2011; Everett and McLeod Jr, 2007). In particular, requirements-driven testing is conducted to validate systems before deployment (Dustin et al., 2009). This is especially critical for systems that must demonstrate functional safety, such as automotive systems. In such contexts, it is also common to specify test cases in natural language (following templates) to facilitate their analysis and traceability to requirements, before manually writing the corresponding test code.
@@
However, due to the increasing complexity of systems and their continuous change by large teams, test suites are hard to maintain, tend to grow, and are often redundant (i.e., some of the test cases are likely to detect the same faults), thus leading to waste of time and resources (Rothermel et al., 2002). Further, in practice, testers have limited resources and time to test new system updates. To address this, given a test budget, test suite minimization (TSM) is proposed to prune test cases that are most likely to be redundant, while satisfying coverage criteria and maximizing fault detection (Harrold et al., 1993; von Ronne, 1999), thus significantly reducing testing cost and time while limiting the impact on system quality or compliance.
@@
While several approaches have been proposed for TSM (Noemmer and Haas, 2019; Miranda and Bertolino, 2017; Coviello et al., 2018; Liu et al., 2011; Viggiato et al., 2022; Cruciani et al., 2019; Zhang et al., 2019; Hemmati et al., 2013; Pan et al., 2023, 2024; Marchetto et al., 2017; Xia et al., 2021; Anwar et al., 2019; Yoo and Harman, 2010), most existing techniques rely on code and structural coverage criteria (e.g., statement or branch coverage). These techniques do not address the unique challenges of minimizing test suites driven by system requirements, with test cases often specified in natural language. Moreover, to the best of our knowledge, no existing solution can simultaneously maintain full requirement coverage while adhering to a fixed minimization budget, a common situation in industrial practice due to limited time and resources.
@@
To bridge this gap, we propose RTM (Requirements coverage-guided Test suite Minimization), a novel approach specifically designed for requirement-based testing. RTM aims to reduce test suite redundancy (in terms of fault detection) while ensuring full requirement coverage, making it particularly well-suited to the practical needs of our industrial partners and to critical systems that must demonstrate functional safety in general. RTM takes requirement-based test cases–composed of natural language test steps–as input and computes pairwise similarity between them. Specifically, we investigate three different preprocessing methods to normalize the test steps, followed by seven distinct text embedding techniques–both sentence-level and word-level–to transform the textual data into vector representations: Term Frequency–Inverse Document Frequency (TF-IDF) (Salton and Buckley, 1988), Universal Sentence Encoder (USE) (Cer et al., 2018), LongT5 (Guo et al., 2022), Amazon Titan Text Embedding V2 (AWS, 2024), Word2Vec (Mikolov et al., 2013), GloVe (Pennington et al., 2014), FastText (Bojanowski et al., 2017). We then measure the similarity between test cases using three different distance metrics: cosine similarity and Euclidean distance for sentence-level embeddings, and Word Mover’s Distance (WMD) (Kusner et al., 2015) for word-level embeddings. Based on the computed similarities, we employ a Genetic Algorithm (GA) (Luke, 2013) to search for the optimal test suite that minimizes similarity, adheres to a fixed budget, and ensures full requirement coverage. To help the GA efficiently find valid subsets, we devise three initialization strategies that generate an initial population of subsets satisfying both coverage and budget constraints.
@@
RTM was evaluated on an industrial automotive dataset consisting of $736$ system test cases across seven test runs, covering $54$ requirements. The results demonstrate that RTM outperforms the baselines in terms of fault detection rate (FDR) under various minimization budgets. Moreover, we explore the impact of test suite redundancy on FDR. Our findings show that: ① RTM consistently outperforms all baseline techniques in terms of FDR across different test suite redundancy levels; ② The level of redundancy has a strong effect on the achieved FDR, for all TSM techniques; ③ Especially at lower minimization budgets and for highly redundant test suites, there is still significant room for improvement between the theoretical maximum FDR and that of RTM. ④ The runtime of RTM scales approximately linearly with test suite size, with small (i.e., of no practical consequence) absolute differences compared to the baselines. Given its consistently higher FDR performance across various test suites with different redundancy levels, RTM is therefore a better choice in many practical contexts.
@@
  • To foster reproducibility and facilitate further research, we have open-sourced our replication package, including source code, input data, and results for all research questions that allow the reproducibility of the experiments, which is publicly available at https://github.com/rongqipan/RTM. Industrial requirements and test cases are confidential.
@@ 2.1. Test Cases based on Automotive Requirements
Like other critical systems, modern automotive software systems are expected to comply with functional safety standards such as ISO 26262 (Siegl et al., 2010). To ensure compliance, system engineers derive and maintain extensive test suites based on system requirements, which are executed to validate system behavior against expected requirements. Typically, test cases are specified in natural language, following templates, and then manually converted into test code and implemented by experts.
@@
$2.$ Test Suite Minimization
Test suite minimization (TSM) aims to reduce the test suite size while ensuring key testing objectives, such as requirement coverage, fault-detection capability, and test-case diversity (Harrold et al., 1993; von Ronne, 1999). Given that test suite diversity is reported to be positively correlated with fault detection capability (Hemmati et al., 2013), minimizing test case similarity can enhance the fault detection capability of the resulting test suite. When appropriately applied, TSM can significantly reduce testing costs and time while minimizing its impact on system quality or compliance.
Let $\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ be the set of system requirements for an automotive system, and let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{m}\}$ be the corresponding test suite. Each test case $t_{i}\in\mathcal{T}$ covers a subset of requirements denoted by $Cover(t_{i})\subseteq\mathcal{R}$. Note that in our dataset, each test case only covers one requirement, whereas many test cases can cover each requirement. This is because each test case is specifically designed to validate a scenario in which the requirement is expected to hold. Therefore, in our context, guaranteeing full requirement coverage requires at least as many test cases as requirements in the test suites. Moreover, covering all requirements is a must in automotive testing, to comply with functional safety standards (Siegl et al., 2010), which are adhered to by our industry partner. Let $Sim(t_{i},t_{j}):\mathcal{T}\times\mathcal{T}\rightarrow[0,1]$ denote the similarity between test cases $t_{i}$ and $t_{j}$.
@@ 3. Approach
This section introduces RTM, a TSM approach that satisfies the three criteria outlined in Section $2.$ It targets test cases derived from system requirements for validation. In short, it leverages a text embedding technique combined with a distance function to compute test case similarity, then employs evolutionary search to minimize the test suite. The primary objective of RTM is to find the optimal $\mathcal{T}^{\prime}$ that maximizes fault detection capability by minimizing the similarity values between the test cases while reducing the testing cost, measured in terms of test suite size, as well as ensuring $100\%$ requirement coverage.
@@
Figure 2. Overall framework of RTM.
@@ 3.2. Test Case Representation
For word-level embeddings (i.e., Word2Vec, GloVe, and FastText), we use WMD (Kusner et al., 2015), which formulates the similarity between two test cases as an optimal transport problem. Specifically, WMD computes the minimum cumulative cost required to move the probability mass of all the words in one test case to those in another in the embedding space. The WMD is defined as:
@@
This optimization problem is equivalent to the Earth Mover’s Distance (Monge, 1781; Wolsey and Nemhauser, 1999; Rubner et al., 1998) and can be solved via linear programming, with a computational complexity of approximately $O(n^{3}\log n)$ for $n$ unique words. Although the computation can be expensive when there are many words, WMD is well-suited to word-level embeddings because pairwise distances between the word embeddings approximate semantic dissimilarity, allowing the transport cost to reflect meaning differences between test cases (Kusner et al., 2015).
In order to bound the WMD between 0 and 11, and convert it from a distance function to a similarity measure, we normalize it as follows:
@@
For sentence-level embeddings (i.e., TF-IDF, USE, LongT5, and Amazon Titan Embedding V2), we investigate both cosine similarity and Euclidean distance, which have been widely used in machine learning, information retrieval, and NLP tasks to measure similarity or dissimilarity between vectors (Gomaa and Fahmy, 2013; Huang et al., 2008; Li and Han, 2013; Salton and Buckley, 1988). However, they operate differently and are suited to different use cases: cosine similarity measures the angle between two vectors, ignoring magnitude, whereas Euclidean distance measures the straight-line distance between two vectors. By applying different types of similarity metrics, we aim to explore how different vector representations and distance functions affect the ability of RTM to detect redundant test cases.
@@ 3.4. Evolutionary Search for Test Suite Minimization
As discussed in Section 2.2, the TSM in our context is an NP-hard problem. Therefore, to efficiently find a near-optimal, feasible solution, we apply a GA–the most widely used evolutionary search algorithm in software testing and the approach adopted by the state-of-the-art (SOTA) TSM work (i.e., ATM (Pan et al., 2023) and LTM (Pan et al., 2024))–which has demonstrated high effectiveness and efficiency for TSM. To tailor GA to our problem, we define the individuals (minimized test suites), the initialization strategy, selection, crossover, and mutation operators, and the fitness function that guides the search. The details are given below.
@@ 3.4.3. Search Process
After initializing a population with valid individuals (subsets), GA (Golberg, 1989) iteratively evolves this population using selection, crossover, and mutation operators towards better individuals with lower fitness values. In each generation, offspring are created by recombining and mutating the genes (i.e., selected test cases) of existing individuals. Each individual is evaluated by computing its fitness, and those with lower fitness are selected for the next generation. The evolutionary process is repeated until convergence, defined as a decrease in the fitness score of less than $0.0025$ across generations. The final output is the best subset identified during the search, which satisfies all constraints and minimizes the redundancy among selected test cases, indicated by the lowest fitness value.
@@
  • Crossover: We used the customized crossover operator (Blank and Deb, 2020) that ensures the offspring maintains a fixed subset size. It first includes all test cases shared by both parent subsets, then adds test cases from either parent subset until the minimization budget is met.
  • Mutation: We used an inversion mutation operator (Blank and Deb, 2020) that randomly selects a segment of the subset and reverses its order. Since a subset is represented as a binary vector, this approach preserves the subset size.
@@
We follow the published guidelines  [^1] https://www.obitko.com/tutorials/genetic-algorithms/recommendations.php used in prior work (Pan et al., 2024) for setting GA hyperparameters. Specifically, we use a population size of $100$, a mutation rate of $0.01$, and a crossover rate of $0.90$. For each generation, the size of each solution (i.e., subset) remains fixed at the minimization budget.
@@ 2. RQ2:
    This RQ aims to compare the performance of RTM with two baseline approaches, namely Random Minimization and FAST-R (Cruciani et al., 2019), under various minimization budgets (i.e., 10%, 20%, …, 90%).